Give an example of infinite ring of characteristic 2


1
Expert's answer
2022-01-31T15:49:44-0500

Consider the infinite ring "(\mathbb Z_2,+,\cdot)^{\omega}=\{(a_0,a_1,a_2,...,a_n,...)|\ a_i\in\Z_2\}," where "\Z_2=\{0,1\}."

Taking into account that

"2\cdot(a_0,a_1,a_2,...,a_n,...)=(2a_0,2a_1,2a_2,...,2a_n,...)=(0,0,0,...,0,...)"

for any "(a_0,a_1,a_2,...,a_n,...)\in(\mathbb Z_2)^{\omega}",

we conclude that this ring is of characteristic 2.